For a given map, two-thirds of an inch represents 80 miles. Two cities are 200 miles apart. How far apart are they on the map?
Answer by checkley79(3341)

80/200=(2/3)/X
80X=200(2/3)
80X=400/3
80X*3=400
240X=400
X=400/240
X=1 2/3 INCHES.
PROOF:
80/200=(2/3)/(5/3)
80(5/3)=200(2/3)
400/3=400/3